Circular No. 6767

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1997 V1 Jeff Larsen reports his discovery of a comet with the 0.91-m Spacewatch

telescope at Kitt Peak. The following observations have been reported:

1997 UT  α_{2000}  δ_{2000}  m_{1} Observer
Nov. 3.22743 0^{h}14^{m}11.^{s}99 +12^{o}28′59′′.4  Larsen
3.24864 0 14 11.70 +12 28 51.3 16.6  _{”}
3.27007 0 14 11.42 +12 28 43.4 16.6  _{”}
3.29130 0 14 11.11 +12 28 35.2  _{”}
3.29561 0 14 11.05 +12 28 33.5 16.6  _{”}
3.29981 0 14 10.99 +12 28 32.0 16.6  _{”}
3.36761 0 14 10.09 +12 28 06.0 16.6  _{”}
3.37237 0 14 10.03 +12 28 04.4 16.7  _{”}
3.37636 0 14 09.97 +12 28 02.8 16.7  _{”}
3.96399 0 14 03.10 +12 24 22.5 16.0 Tichý
3.96612 0 14 03.04 +12 24 21.3  _{”}
3.96756 0 14 03.02 +12 24 20.6  _{”}
3.97039 0 14 02.98 +12 24 20.2  _{”}

J. Larsen (Kitt Peak). Measurer J. V. Scotti. Coma diameter 14′′; tail

extends 0′.89 in p.a. 258^{o}; m_{2} = 19.2–19.7. M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector + CCD. Measurer

Tichý. Coma diameter 11′′; no tail. Not very good seeing.

EF PEGASI J. A. Mattei, AAVSO, reports that T. Burrows (Novato, CA) and P.

Schmeer (Bischmisheim, Germany) independently detected a superoutburst of this SU UMa-type dwarf nova, the first since early 1995 (IAUC 6125), when the star reached maximum at mag 10.7 by Jan. 12 and remained bright for three weeks. Visual magnitude estimates: Oct. 31.842 UT, [14.4 (E. Broens, Mol, Belgium); Nov. 1.93, [13.0 (P. Schmeer, Bischmisheim, Germany); 2.187, 11.4 (T. Burrows, Novato, CA); 2.713, 11.1 (Schmeer); 2.835, 10.7 (Schmeer); 3.077, 11.2 (G. Hanson, Cave Creek, AZ); 3.160, 10.9 (R. Royer, Lakewood, CA); 3.208, 11.3 (Hanson); 3.779, 11.2 (Broens); 3.795, 11.3 (E. Muyllaert, Oostende, Belgium). A star of mag 12.4 is located only ∼ 5′′ from EF Peg, at p.a. 305^{o}.

1997 November 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green